OpenStudy (anonymous):

X^2-488x+4780 find the factors

jimthompson5910 (jim_thompson5910):

you could do this by trial and error (to find two numbers that multiply to 4780 and add to -488) or you could use the quadratic formula

jimthompson5910 (jim_thompson5910):

in this case a = 1, b = -488 and c = 4780

OpenStudy (anonymous):

Do u know any numbers

jimthompson5910 (jim_thompson5910):

\[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(-488)\pm\sqrt{(-488)^2-4(1)(4780)}}{2(1)}\] \[\Large x = \frac{488\pm\sqrt{238144-(19120)}}{2}\] \[\Large x = \frac{488\pm\sqrt{219024}}{2}\] I'll let you finish

OpenStudy (anonymous):

I got x=10 and x=478

OpenStudy (anonymous):

Thanks for Your help it was really helpful and useful I understand this now

jimthompson5910 (jim_thompson5910):

sry i was afk the two solutions are x=10 or x=478, so you are correct this means that the two numbers are -478 and -10 (you flip the signs) so -478 + (-10) = -488 and -478 * (-10) = 4780 which means x^2-488x+4780 factors to (x - 478)(x - 10)

OpenStudy (anonymous):

Yes